the sbc bookstore sold 348 notebooks during the first day of school, some at P 22.50 each and the rest at P 34.20 each. the total amount taken in for the notebooks was taken in for the notebooks was P 9081.90. How many of each kind were sold?

Answer by checkley71(8403)

X+Y=348 X=348-Y
22.50X+34.2*Y=9081.90
22.50(348-Y)+34.2Y=9081.90
7,830-22.5Y+34.2Y=9081.90
11.7Y=9081.90-7,830
11.7Y=1,251.9
Y=1,251.9/11.7
Y=107 $34.20 BOOKS WERE SOLD.
348-107=241 $22.50 BOOKS WERE SOLD.
PROOF:
22.50*241+34.2*107=9081.9
55422.5+3659.4=9081.9
9081.9=9081.9